OpenStudy (anonymous):

Solve logx + 5 36 = 2 Choose one answer. a. x = 1 b. x = 11 c. x = 13 d. x = 23

OpenStudy (anonymous):

\[\log_{x+5} (36)=2\]

OpenStudy (anonymous):

\[\log_{x+5}(36)=\ln (6^{2})\div \ln (x+5)\]

OpenStudy (anonymous):

so the équation become \[\ln (6)=\ln (x+5)\]

OpenStudy (anonymous):

ok what do i do next

OpenStudy (anonymous):

I think its x=1 but I am not positive it that right

OpenStudy (anonymous):

So you can easily conclude that x+5 = 6 x = 1 So the answer is (a)

OpenStudy (anonymous):

kk thanks so much for all your help!

OpenStudy (anonymous):

to be sure you can replace x by 1.

OpenStudy (anonymous):

\[\log_{6} (36)=\ln(6^{2})\div \ln (6)=2\ln (6)\div \ln (6)=2\]